Consider a deep pit solid waste disposal facility constructed in Iim thick horizontal saturated clay layer Underneath the clay layer, there is a sand aquifer with constant thickness (10m) In the waste Leachate, two chemicals, Lead (Pb) and Chloride (CI) are dissolved in high concentration Two monitoring wells are installed at the upstream side (Well_1) and downstream side (Well 2).

Datum line to define the elevation bead: Ground surface z= 0 m

Leachate level in the pit: Ground surface 

Hydraulic conductivity of clay= 10^-9 m/s

Hydraulic conductivity of sand layer = 10^-9m/s

 Porosity of clay= 0.5

 density of soil grain= 2.65 gr/cm^3

Effective diffusion coefficient of clay: 2×10^-9 m^2/s

water depths of Well_1= 5m and Well 2= 9m 

In the sand layer horizontal flow can be approximated from Well_1 to Well 2. 

One dimensional vertical flow can be assumed in the 3m thick clay

 beneath the pit

(1)How much are the void ratio (e) and moisture content (w) of the clay?




1
Expert's answer
2021-04-01T01:14:53-0400

e=vv/vs

vs=0.5/2650 = 0.0001886

vv=0.00003-0.0001886

=0.0001586

=0.0001586/0.0001886

e=0.841

m=mw/ms

9-5/5

0.8